FEL Upgrade Project Weekly Brief
August 6-10, 2001

This was the last week of the scheduled Summer User run on the FEL which was devoted to a continuation of Bob Austin's experiments on vibrational energy transport in proteins.  Next week we have scheduled additional measurements of terahertz radiation with the help of our colleagues from Brookhaven.

Papers were prepared for presentation at the International FEL Conference which starts on Aug. 19th in Darmstadt.
Presentations that were reviewed this week included:

George Neil's overview of light source development in the last decade and a talk on second harmonic generation in the IR Demo; Steve Benson's review of lesson's learned with the IR Demo FEL; Lia Merminga's analysis of RF physics and beam current limits in energy recovered linacs; and Ed Gillman's review of his user experiments with scanning near field microscopy. When the accompanying manuscripts have been completed and reviewed they will be made available on our server.

We M. Shinn and G. Neil attended a "Lessons Learned" meeting at JTO in Washington on Monday.  Some initial results of the laser materials testing were presented.  There is much interest in the results and several modifications to the test approach for September were discussed and agreed upon.  A small action item list for fall test preparation was released.

As part of our normal end of year financial review, we continued our analysis of the projected FY02 budget, manpower and resource allocations. Next week (Aug.17), DOE will be reviewing Jefferson Lab's Institutional Plan.

WBS 4 (Injector):
The anode support plate drawing was signed off and submitted to the shop for fabrication.  The anode plate polishing was completed to 1 um.  NEG pumps for the gun chamber were ordered.
WBS 5 (SRF):
o Qualification of JL008 was completed this week with final gradient of 16.3 MV/m at Q value of 6.5e9. This cavity will now be completed with addition of end groups and helium vessel.
o Helium vessel / endgroup alignment tooling is being modified to accomidate the latest revisions to the flanges.
o The new FPC flange testing was completed with successful cryo-cycles to 2 Kelvin liquid helium and multiple tests. This completes all the flange tests scheduled.
o Higher Order Mode testing is continuing on the final positions of the coupler probes.
o A second vertical test cavity stand was completed and is undergoing leakchecking. This second stand will increase the number of cavity tests allowing for faster turnaround in this qualification stage.
WBS 6 (RF):
Zone 3 - Received 4 RF Control Modules and 4 partial RF Control modules.  The wiring of the zone is 95 % complete.  EPICS is connected and communicating with the RF Control Modules.  In 1.5 weeks the current will be over.  This will allow the vault to be opened for additional cables and test equipment to be connected and the 480 VAC to be connected to the klystrons' HVPS.

Zone 4 - No progress.  In appears some items have been removed.  We will arrange to have those reinstalled.

Quarter Klystrons - September 12th has been set for the factory tests at CPI.  The first cart for the 100 kW klystrons is still in the Machine Shop.

Quarter HVPS - The SCR Controller to ramp the HVPS on and off has been ordered.  The rectifier assemblies are also on order.

Gun HVPS - When the vault can be opened, the multiplier and monitoring stacks will be assembled for the mechanical engineer to measure for mounting them in the new HVPS Tank.  A proposed schedule has been offered for the design, modeling, drafting, and construction of the tank, internal hanger assembly, transmission line, and the support assembly for the HVPS driver electronics.  The proposed schedule runs until mid-June, but shows only 1 engineer and 1 designer working on the project where 2 of each could greatly reduce the time involved.  We are man-power limited.

WBS 8 (Instrumentation):
Testing continues on the VME timing/delay board.  The new 32 channel sample and hold buffer board is ready for prototyping.  Boards have been ordered.  Prototype boards for the new beamviewer camera/lamp board power supplies have also been ordered and a bracket designed for mounting the power supply and air solenoid to the beamviewer housing.

The "Rabbit" programmable TCP/IP controller board was successfully used for the first time.  It was programmed to interface to LabView for Bob Austin's experiment.  The Rabbit controlled the direction of the sevo interferometer.

Trim rack specifications are nearly complete. The racks will go out for bid for 4 racks; 128 channels with an option for an additional 128 channel (if $ is available). The plan is to use the new DSP/4-quadrant trims to control the iron core magnets and to use the existing trim channels on the air core magnets. The RFP will go out for complete rack assemblies.

The first of the new Ion Pump Power Supplies is expected to be delivered from the assembler soon for testing.

The BAIR web and the task system have received several updates.  An "UPGRADE" section was added to BAIR web for information on the 10KW upgrade.  Anyone with information to post can contact Mike Aston directly via a link on the site.  The task system has been updated to allow searching for specific tasks by system. Personnel wanting information on a particular system can enter that system name, and all tasks associated with that system are available.

Drawings completed this week:  F0060 - VME Timing Main Board. Several others are nearing completion.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  We posted a copy of the solid model (drafting file) on a Web Site for access by bidders to use in the manufacturing planning.
Injector Dipoles (DU/DV)
o  The remainder of the detail drawings of the Small Injector Dipoles (GV) were received from DULY and went through another check cycle and were sent back to DULY for minor corrections.  The are completing work on the first round of rework on the Assembly Drawing.
Arc Dipoles (GY, GX, GQ)
o  In magnetic modeling of the 180 Dipole (GY), AES is nearly complete with the final adjustments, the end face position and the angle of the path length coil slot to bring the field integrals at any radius within the good field region into 180 degrees of effective length.
o  Design continued on correcting the drawings of the GY (they were checked several months ago) and bringing them up to the latest magnetic model.
3 inch quad (QX)
o  The steel for cores is in shipment to New England Technicoil for delivery early next week.
o  Magnet Enterprises International is continues perfecting their potting technique on the first article coil.  Al Guerra and Jeff Dale will inspect the first article next week.
Trim Quad (QT)
o  The 2D detail drawings are being checked.
Sextupole (SF)
o  Drawing layout continued.
Octupole (OT)
o  We reached agreement on the scope of the Task Order with DULY Research to do magnetic modeling and find out how difficult it will be to obtain a 1% magnet from a configuration that is "all ends".
Corrector Dipoles (DB, DJ)
o  We are planning to visit Milhous Co. next week to assess progress in manufacturing the corrector coil order.
o  We laid out a design of the new mounting bracket that will hold the DB/DJ correctors over the new larger diameter beam position detectors.
Beam Line and Vacuum
o  Work continued on the Interface Control Document for the task definition for the Arc and Optical Chicane chambers.  Work concentrated on the diagnostics, flanges and magnets at the center of the Optical Chicane.
o  The bid package drawing set for the girders in the return leg continues in detailing.
o  We worked on bringing the Fast Track Schedule up to date and filling in detail where there were general tasks.

WBS 10 (Wiggler):
Efforts this week focussed on getting the measurement stand calibrated.  An offset in the longitudinal position of 400 microns was found.  Efforts are now underway to fix this problem before measurements commence.

WBS 11 (Optics):
Design progress on the optical cavity chambers continues, with continued refinement of the outcoupler mirror holder fixturing.  This fixturing is also consistent with the high reflector mounting.  The vertical motion design is also well along, with the actuation, leadscrew, and pillow block defined.  Thus, we are close to being able to place orders for some of the longer lead time components.
The RFP for the new drive laser was sent out.  The RFP for the ultrafast laser (to be used as a diagnostic) is being drafted.

Support for operations/experimenters:
Changed mirrors and diagnostics for operations at 6 microns.  Did further work on autocorrelator, which has been going slowly because of the low duty factor.  Provided support for the primary user (R. Austin - Princeton Univ.).  Continued calibration of the IR camera on the mirror test stand, so we can determine the losses of our mirrors.  Prepared for measurements we'll do tomorrow (8/10).  Performed tests of the epoxy we will use to affix mirrors to the deformable mirror assembly.  Measured the phase noise of the drive laser.

On Monday, G. Neil and I traveled to Arlington, VA, for a "lessons learned" meeting with officials from the JTO, collaborators from the NRL, and observers from Lincoln Labs.  We also presented the data analyses from one target material.
We ran the FEL all week for users.  The primary activity by the users from Princeton and the Netherlands was measurement of energy transfer mechanisms in biological molecules, specifically myoglobin.  The FEL operated normally in pulsed mode at 5.9 to 6 microns.  The final day of running was devoted to measurements of high reflectivity mirror coating absorption at 6 microns. These tests were underway as this report was written.  Tests next week are planned for beam breakup and higher order mode measurements, Terahertz wave generation and study, and a continuation of polymer surface modification.  This will conclude our extremely productive summer run.

Privacy and Security Notice